OpenStudy (anonymous):

Which of the following is an equivalent representation of 5^-4? Choices: 1/256 1/64 1/16 1/625

OpenStudy (anonymous):

Any negative exponent means it's a reciprocal, i.e.: \[ a^{-b}=\frac{1}{a^b} \]So: \[ 5^{-4}=\frac{1}{5^4} \]Try it, now.